# Interplay between local and global order controls bifurcations in nonreciprocal systems without a conservation law

### v1

@@ Abstract
Hence, the Hopf bifurcation occurs at $J=1/z$ and $K\neq 0$, such that ${\rm Re}(\lambda^{\rm MF}_{\pm})=0$ and ${\rm Im}(\lambda^{\rm MF}_{\pm})\neq 0$. This corresponds to a straight vertical line in the $(J,K)$-plane, as shown in Fig. A1a, which is unphysical for short range interactions. Notably, in the MF approximation we do not observe a critical value for short-range order, which is present in the more accurate BG approximation.
Figure A1: Phase diagram for global order $m^{\mu}_{s}$ obtained with the mean-field (MF) approximation (a) and the monomer approximation (b) for the perfect nonreciprocal setting with $J_{a}=J_{b}=J$ and $K_{a}=-K_{b}=K$. In the MF approximation the Hopf bifurcation (blue line) is set by $J=1/z$. In the monomer approximation, the Hopf bifurcations (blue line) is set by a critical short-range order $q_{s}=1/z$, which is more similar to the Bethe-Guggenheim approximation where $q_{s}=1/(z-1)$ (see Fig. 2d).
@@
and $\mathcal{P}^{b}_{l,n}(t)$ is obtained by replacing $m^{a}(t)$ with $m^{b}(t)$ in Eq. (B2). Since $\mathcal{P}^{a}_{l,n}(t)$ is independent of the short-range order $q^{\mu\nu}(t)$, this implies that $q^{\mu\nu}(t)$ is slaved by $m^{\mu}(t)$, however, $q^{\mu\nu}(t)\neq m^{\mu}(t)m^{\nu}(t)$. Performing a linear stability analysis around the trivial steady-state $m^{\mu}_{s}=0$ for $J_{a}=J_{b}=J$ and $K_{a}=-K_{b}=K$, we obtain the linear stability equation $\tau d\delta\mathbf{m}(t)/dt=\mathbf{\hat{M}}(J,K)\delta\mathbf{m}(t)$ where the elements of $\mathbf{\hat{M}}(J,K)$ are given by
@@
Interestingly, Eq. (B3) can be expressed in terms of the steady-state value of the short-range order in the monomer approximation, i.e., $\hat{M}_{aa}(J,K)=zq_{s}(J,K)-1$. Therefore, the eigenvalues of $\mathbf{\hat{M}}(J,K)$ are neatly given by
@@
Herefrom it follows that the Hopf bifurcation occurs at $q_{s}=1/z$ and $K\neq 0$, such that ${\rm Re}(\hat{\lambda}_{\pm})=0$ and ${\rm Im}(\hat{\lambda}_{\pm})\neq 0$. Hence, as within the BG approximation, we also find the existence of a critical short-range order in the monomer approximation. Contrary to the MF approximation, the Hopf line is not a straight vertical line in the $(J,K)$-plane, as shown in Fig. A1b.
@@
In this Supplemental Material, we provide derivations and further details about the results shown in the Letter. Furthermore, in Sec. S6 we prove that static long-range order in the square-lattice Ising model persists under small perturbations in the nonreciprocal coupling $K$, which invalidates a recent hypothesis [1].
S1 Setup: Dual lattice topology and interactions
@@
where $\Delta E^{\mu}_{i}$ is the change in energy on the $\mu=a\lor b$ lattice after flipping spin $\sigma^{\mu}_{i}\rightarrow-\sigma^{\mu}_{i}$, and $\tau\equiv 2N$ is an intrinsic timescale to attempt a spin-flip. According to Eq. (S1), the change in energy after such a spin-flip is given by
@@
where $\langle f(t)\rangle\equiv\sum_{\boldsymbol{\sigma}}P(\boldsymbol{\sigma};t)f(\boldsymbol{\sigma})$. Eqs. (S5)-(S6) are exact (but not yet closed) equations for the first two moments evolving under Glauber dynamics, and will serve as our starting point to derive equations for the long- and short-range order parameters.
S3 Differential equations for the long- and short-range order
Recall from the Letter that the long- and short-range order are defined as
@@
where $N^{\mu}_{+}$ is the total number of up spins on the $\mu$ lattice, and similarly $N^{\mu}_{-}$ the total number of down spins. The same reasoning applies to the probability $\mathcal{P}^{b}_{l,n}$. Note that $N^{\mu}_{\pm}$ is generally time-dependent, but for simplicity we omit the subscript $t$. To relate $N^{\mu}_{\pm}$ to the long-range order we use the following relations (omitting the time dependence subscript $t$)
@@
Hence, in the monomer approximation, we find that the probability $\mathcal{P}^{\mu}_{l,n}(t)$ only explicitly depends on the long-range order $\mathbf{m}(t)$, and not on the short-range order $\mathbf{q}(t)\equiv(q^{aa}(t),q^{bb}(t),q^{ab}(t))$. Upon inserting Eq. (S27) into Eqs. (S19)-(S21) we obtain a closed set of five non-linear differential equations, where, notably, the dynamical equations for $\mathbf{q}(t)$ are slaved by the solutions for $\mathbf{m}(t)$, as the r.h.s. of Eqs. (S19)-(S21) do not depend on $\mathbf{q}(t)$.
@@
where, for example, $N^{ab}_{+-}$ is the total number of nearest-neighbor spin pairs with an up spin on the $a$ lattice and a down spin on the $b$ lattice. To relate $N^{ab}_{+-}$ and the other spin pair numbers to the short- and long-range order, we make use of the following exact relations for periodic lattices
@@
in combination with Eq. (S25). Furthermore, we use the definition of short-range order (see Eqs. (S8)-(S9)) to write
@@
where $\mathcal{C}^{z}_{l}$ is given in Eq. (S29). Expressions (S43)-(S44) are given in the main Letter in Eq. (6). The expressions for $\mathcal{P}^{b\pm}_{l,n}(t)$ are similar to Eqs. (S43)-(S44) upon interchanging $m^{a}(t)$ and $m^{b}(t)$ and replacing $q^{aa}(t)$ with $q^{bb}(t)$. The difference with the “monomer” approximation (S27) and (S28) is that in the BG approximation the probability also depends explicitly on the short-range order $\mathbf{q}(t)$. Inserting Eqs. (S43)-(S44) into Eq. (S31), and finally into Eqs. (S19)-(S21), leads to a closure of the five coupled non-linear differential equations.
S4 Linear stability analysis
@@
Since in the MF approximation the short-range order is slaved by the long-range order, it suffices to consider the linear stability analysis for the long-range order. Let us consider a small perturbation of the long-range order around the steady-state value $\textbf{m}_{\rm s}$, i.e., $\mathbf{m}(t)=\mathbf{m}_{\rm s}+\mathbf{\delta m}(t)$. The most interesting steady-state value to consider is given by the disordered state $\mathbf{m}_{\rm s}=0$, in which case the probability (S27)-(S28) can be expanded as
@@
Note that $\hat{M}_{\mu\mu}$ can be written in terms of the the steady-state solution for the short-range order given by Eq. (S20), i.e., $\hat{M}_{\mu\mu}=zq^{\mu\mu}_{s}-1$, which we will make use of in the remaining calculation. Furthermore, the off-diagonals are related to the steady-state of the short-range order between the lattices, i.e., $\hat{M}_{ab}+\hat{M}_{ba}=2q^{ab}_{s}$. Hence, a perturbation of the long-range order couples to the steady-state value of the short-range order. The solution of Eq. (S47) can be written in terms of an eigenmode expansion
@@
For fixed $z$, both equations can be solved explicitly to obtain expressions for the critical values of the parameters $J_{\mu}$ and $K_{\mu}$ on the line of Hopf bifurcations. Note that Eq. (S56) sets a direct constraint on the steady-state short-range order values.
@@
where the latter equation is directly satisfied for $K\neq 0$ (a proof is given in S4.2.2.). Hence, also in the “monomer” expression we find a critical value for the short-range order, which is given by $q_{\rm crit}\equiv 1/z$. Upon inserting Eq. (S48) into Eq. (S58), one can obtain an explicit result for the Hopf line in the $(J,K)-$plane.
@@
In the BG approximation, we can no longer neglect the short-range order for the linear stability analysis. Hence, we consider a small perturbation of the long- and short-range order around their respective steady-state values
@@
S4.2.1 Steady-state short-range order
The trivial steady state is given by the disordered state with $m^{\mu}_{s}=0$ and $q^{ab}_{s}=0$. To solve for the steady state of the short-range order, denoted as $q^{aa}_{s}=q^{bb}_{s}\equiv q_{s}(J,K)$, we need to solve
@@
S4.2.2 Linear stability of long-range order
@@
S4.2.3 Linear stability of short-range order
@@
S5 Divergence of the residence time upon approaching the SNIPER bifurcation
@@
S6 Proof of existence of static long-range order on the square lattice
In [1] it is claimed that for the perfect nonreciprocal setting with $J_{a}=J_{b}=J$ and $K_{a}=-K_{b}=K$, any nonzero nonreciprocal coupling $K\neq 0$ destroys the static long-range order in the two-dimensional square-lattice Ising model. The authors’ reasoning behind this claim goes as follows: Consider the low temperature limit, and suppose all spins for the two lattices are either up or down. Then, the total “energy” in both lattices can be written as
$E^{a}$ $=-J\sum_{\langle ij\rangle}\sigma^{a}_{i}\sigma^{a}_{j}-KN,$ (S92)
$E^{b}$ $=-J\sum_{\langle ij\rangle}\sigma^{b}_{i}\sigma^{b}_{j}+KN.$ (S93)
Clearly, lattice $b$ is in a meta-stable state, and therefore tries to flip all spins, thereby transiting to the global minimum. But if this happens, lattice $a$ is in a meta-stable state, and therefore all spins in lattice $a$ try to flip to reach the global minimum. This results in a repetitive cycle of consecutive magnetization reversal in both lattices, thereby destroying the static-order.
One should immediately note, however, that this reasoning is not restricted to the square lattice, but also to any other lattice (incl. the fully-connected mean-field lattice), which would indicate that static long-range order is destroyed for any lattice. This is clearly not the case, as the authors themselves show in [1] for the mean field lattice, and we show in the present Letter also for the Bethe lattice. Furthermore, the above argumentation based on gradient descent in the “energy” landscape does not apply to a non-equilibrium, detailed-balance violating dynamics.
To resolve this obvious inconsistency, we prove that static long-range order in the square lattice nonreciprocal Ising model persists upon small perturbations in the nonreciprocal coupling strength. Thereby, we directly disprove that static-order vanishes in the two-dimensional nonreciprocal Ising model with $K\neq 0$ as stated in [1].
S6.1 Recap of the long-range order
@@
Let us first recall the steady-state long-range order for $K=0$, where we have two independent Ising systems on lattice $a$ and $b$, respectively. Let $P_{0,\rm s}(\boldsymbol{\sigma})$ and $w_{0,i}^{\mu}(\sigma^{\mu}_{i})$ be the steady-state probability and transition rate when $K=0$, respectively. Note that $\sum_{\{\boldsymbol{\sigma}\}}P_{0,\rm s}(\boldsymbol{\sigma})=1$. The steady-state long-range order for $K=0$, denoted as $m^{\mu}_{0,\rm s}$, reads [5]
$m^{\mu}_{0,\rm s}\equiv N^{-1}\sum_{\{\boldsymbol{\sigma}\}}\sum_{i=1}^{N}\sigma^{\mu}_{i}P_{0,\rm s}(\boldsymbol{\sigma})=\pm\left(1-\sinh^{-4}{(2J)}\right)^{1/8},$ (S97)
which is nonzero for $J>\ln{(1+\sqrt{2})}/2$.
@@
$\delta w_{i}^{\mu}(\sigma^{\mu}_{i})=(1/2\tau)[1-2\delta_{\mu,b}]\sigma^{a}_{i}\sigma^{b}_{i}\operatorname{sech}^{2}{(J\sum\nolimits_{\langle i|j\rangle}\sigma^{\mu}_{j})},$ (S99)
@@
$\delta m^{\mu}_{\rm s}\equiv N^{-1}\sum_{\{\boldsymbol{\sigma}\}}\sum_{i=1}^{N}\sigma^{\mu}_{i}\delta P_{\rm s}(\boldsymbol{\sigma}).$ (S101)
@@
S6.4 Bound on perturbations of long-range order
@@
$\delta m^{\mu}_{\rm s}$ $\leq|N^{-1}\sum_{\{\boldsymbol{\sigma}\}}\sum_{i=1}^{N}\sigma^{\mu}_{i}\delta P_{\rm s}(\boldsymbol{\sigma})|$
$\leq N^{-1}\sum_{\{\boldsymbol{\sigma}\}}\sum_{i=1}^{N}|\sigma^{\mu}_{i}\delta P_{\rm s}(\boldsymbol{\sigma})|$
$=N^{-1}\sum_{\{\boldsymbol{\sigma}\}}\sum_{i=1}^{N}|\sigma^{\mu}_{i}||\delta P_{\rm s}(\boldsymbol{\sigma})|$
=$=\sum_{\{\boldsymbol{\sigma}\}}|\delta P_{\rm s}(\boldsymbol{\sigma})|.$ (S102)
@@
$\delta m^{\mu}_{\rm s}$ $\geq-|N^{-1}\sum_{\{\boldsymbol{\sigma}\}}\sum_{i=1}^{N}\sigma^{\mu}_{i}\delta P_{\rm s}(\boldsymbol{\sigma})|$
$\geq-N^{-1}\sum_{\{\boldsymbol{\sigma}\}}\sum_{i=1}^{N}|\sigma^{\mu}_{i}\delta P_{\rm s}(\boldsymbol{\sigma})|$
$=-N^{-1}\sum_{\{\boldsymbol{\sigma}\}}\sum_{i=1}^{N}|\sigma^{\mu}_{i}||\delta P_{\rm s}(\boldsymbol{\sigma})|$
=$=-\sum_{\{\boldsymbol{\sigma}\}}|\delta P_{\rm s}(\boldsymbol{\sigma})|$ (S103)
@@
$|\delta m^{\mu}_{\rm s}|\leq\sum_{\{\boldsymbol{\sigma}\}}|\delta P_{\rm s}(\boldsymbol{\sigma})|.$ (S104)
This is our first main result. We are left with determining an upper bound for $\sum_{\{\boldsymbol{\sigma}\}}|\delta P_{\rm s}(\boldsymbol{\sigma})|$, for which we use the following theorem shown in [6] (Theorem 2.1):
@@
Since it is known that (infinite volume) Glauber dynamics on $\mathbb{Z}^{d}$ is exponentially weakly ergodic (see Theorem 3.3 on page 117 in [7]), we can directly use Theorem 1. To translate the results from Theorem 1 into a bound for $|\delta m^{\mu}_{\rm s}|$, we first want to bound the subordinate norm for the perturbed infinitesimal generator. To do so, note that the perturbation to the transition rates obeys the following bound,
$|K\delta w_{i}^{\mu}(\sigma^{\mu}_{i})|=|(K/2\tau)[1-2\delta_{\mu,b}]\sigma^{a}_{i}\sigma^{b}_{i}\operatorname{sech}^{2}{(J\sum\nolimits_{\langle i|j\rangle}\sigma^{\mu}_{j})}|\leq|K/2\tau|=|K|/4N,$ (S107)
@@
where we used that $|\operatorname{sech}^{2}(x)|\leq 1$ for $x\in\mathbb{R}$ and $\tau\equiv 2N$. These perturbations enter the off-diagonal terms of $\mathbf{\hat{A}}$, so we have established that $\hat{A}_{ij}\leq|K|/4N$ for $i\neq j$. Since we consider single spin-flip dynamics, each row/column in $\mathbf{\hat{A}}$ has $2N$ nonzero entries excluding the diagonal entry, which is equal to $\hat{A}_{ii}=-\sum_{j}\hat{A}_{ij}$. Therefore, the diagonal term can also be bounded by $|\hat{A}_{ii}|\leq 2N\times|K|/4N$. Combining these results, we obtain the following bound for the subordinate norm of the perturbed infinitesimal generator,
$\|\mathbf{\hat{A}}\|\leq 2\times(2N)\times|K|/4N=|K|.$ (S108)
@@
Note, that this result is independent of the system size $N$. Finally, we note that $\|\mathbf{p}_{\rm s}-\mathbf{\hat{p}}_{\rm s}\|$ in Theorem 1 translates into $|K|\sum_{\{\boldsymbol{\sigma}\}}|\delta P_{\rm s}(\boldsymbol{\sigma})|$ in our work. Combining all together, we obtain that there exists $b>0$ and $c>2$ such that for $K\neq 0$
$\sum_{\{\boldsymbol{\sigma}\}}|\delta P_{\rm s}(\boldsymbol{\sigma})|\leq\frac{1+\ln{(c/2)}}{b}<\infty,$ (S109)
@@
and therefore the perturbation is bounded by a finite number independent of $K$ up to first order
$|\delta m^{\mu}_{\rm s}|\leq\frac{1+\ln{(c/2)}}{b}<\infty.$ (S110)
S6.5 Connecting the dots and final remark

### v2

@@ Abstract
Hence, the Hopf bifurcation occurs at $J=1/z$ and $K\neq 0$, such that ${\rm Re}(\lambda^{\rm MF}_{\pm})=0$ and ${\rm Im}(\lambda^{\rm MF}_{\pm})\neq 0$. This corresponds to a straight vertical line in the $(J,K)$-plane, as shown in Fig. A1a, which is unphysical for short range interactions. Notably, in the MF approximation we do not observe a critical value for local order, which is present in the more accurate BG approximation.
Figure A1: Phase diagram for global order $m^{\mu}_{s}$ obtained with the mean-field (MF) approximation (a) and the monomer approximation (b) for the perfect nonreciprocal setting with $J_{a}=J_{b}=J$ and $K_{a}=-K_{b}=K$. In the MF approximation the Hopf bifurcation (blue line) is set by $J=1/z$. In the monomer approximation, the Hopf bifurcations (blue line) is set by a critical local order $q_{s}=1/z$, which is more similar to the Bethe-Guggenheim approximation where $q_{s}=1/(z-1)$ (see Fig. 2d).
@@
and $\mathcal{P}^{b}_{l,n}(t)$ is obtained by replacing $m^{a}(t)$ with $m^{b}(t)$ in Eq. (B2). Since $\mathcal{P}^{a}_{l,n}(t)$ is independent of the local order $q^{\mu\nu}(t)$, this implies that $q^{\mu\nu}(t)$ is slaved by $m^{\mu}(t)$, however, $q^{\mu\nu}(t)\neq m^{\mu}(t)m^{\nu}(t)$. Performing a linear stability analysis around the trivial steady-state $m^{\mu}_{s}=0$ for $J_{a}=J_{b}=J$ and $K_{a}=-K_{b}=K$, we obtain the linear stability equation $\tau d\delta\mathbf{m}(t)/dt=\mathbf{\hat{M}}(J,K)\delta\mathbf{m}(t)$ where the elements of $\mathbf{\hat{M}}(J,K)$ are given by
@@
Interestingly, Eq. (B3) can be expressed in terms of the steady-state value of the local order in the monomer approximation, i.e., $\hat{M}_{aa}(J,K)=zq_{s}(J,K)-1$. Therefore, the eigenvalues of $\mathbf{\hat{M}}(J,K)$ are neatly given by
@@
Herefrom it follows that the Hopf bifurcation occurs at $q_{s}=1/z$ and $K\neq 0$, such that ${\rm Re}(\hat{\lambda}_{\pm})=0$ and ${\rm Im}(\hat{\lambda}_{\pm})\neq 0$. Hence, as within the BG approximation, we also find the existence of a critical local order in the monomer approximation. Contrary to the MF approximation, the Hopf line is not a straight vertical line in the $(J,K)$-plane, as shown in Fig. A1b.
@@
In this Supplemental Material, we provide derivations and further details about the results shown in the Letter. Furthermore, in Sec. S6 we prove that static global order in the square-lattice Ising model persists at least under small perturbations in the nonreciprocal coupling $0<|K|\leq\mathcal{O}(1/N)$ for any finite but arbitrarily large system size $N\in[m,\infty)$ for some sufficiently large $m$.
Appendix S1 Setup: Dual lattice topology and interactions
@@
where $\Delta E^{\mu}_{i}$ is the change in energy on the $\mu=a\lor b$ lattice after flipping spin $\sigma^{\mu}_{i}\rightarrow-\sigma^{\mu}_{i}$, and $\tau$ is an intrinsic timescale to attempt a single spin-flip. In a discrete-time setting the intrinsic flip rate $1/\tau$ is replaced by the probability to flip a single spin, which is $1/2N$ (see also Fig. S2). According to Eq. (S1), the change in energy after such a spin-flip is given by
@@
where $\langle f(t)\rangle\equiv\sum_{\boldsymbol{\sigma}}P(\boldsymbol{\sigma};t)f(\boldsymbol{\sigma})$. Eqs. (S5)-(S6) are exact (but not yet closed) equations for the first two moments evolving under Glauber dynamics, and will serve as our starting point to derive equations for the global and local order parameters.
Appendix S3 Differential equations for the global and local order
Recall from the Letter that the global and local order are defined as
@@
where $N^{\mu}_{+}$ is the total number of up spins on the $\mu$ lattice, and similarly $N^{\mu}_{-}$ the total number of down spins. The same reasoning applies to the probability $\mathcal{P}^{b}_{l,n}$. Note that $N^{\mu}_{\pm}$ is generally time-dependent, but for simplicity we omit the subscript $t$. To relate $N^{\mu}_{\pm}$ to the global order we use the following relations (omitting the time dependence subscript $t$)
@@
Hence, in the monomer approximation, we find that the probability $\mathcal{P}^{\mu}_{l,n}(t)$ only explicitly depends on the global order $\mathbf{m}(t)$, and not on the local order $\mathbf{q}(t)\equiv(q^{aa}(t),q^{bb}(t),q^{ab}(t))$. Upon inserting Eq. (S27) into Eqs. (S19)-(S21) we obtain a closed set of five non-linear differential equations, where, notably, the dynamical equations for $\mathbf{q}(t)$ are slaved by the solutions for $\mathbf{m}(t)$, as the r.h.s. of Eqs. (S19)-(S21) do not depend on $\mathbf{q}(t)$.
@@
where, for example, $N^{ab}_{+-}$ is the total number of nearest-neighbor spin pairs with an up spin on the $a$ lattice and a down spin on the $b$ lattice. To relate $N^{ab}_{+-}$ and the other spin pair numbers to the global and local order, we make use of the following exact relations for periodic lattices
@@
in combination with Eq. (S25). Furthermore, we use the definition of local order (see Eqs. (S8)-(S9)) to write
@@
where $\mathcal{C}^{z}_{l}$ is given in Eq. (S29). Expressions (S43)-(S44) are given in the main Letter in Eq. (6). The expressions for $\mathcal{P}^{b\pm}_{l,n}(t)$ are similar to Eqs. (S43)-(S44) upon interchanging $m^{a}(t)$ and $m^{b}(t)$ and replacing $q^{aa}(t)$ with $q^{bb}(t)$. The difference with the “monomer” approximation (S27) and (S28) is that in the BG approximation the probability also depends explicitly on the local order $\mathbf{q}(t)$. Inserting Eqs. (S43)-(S44) into Eq. (S31), and finally into Eqs. (S19)-(S21), leads to a closure of the five coupled non-linear differential equations.
Appendix S4 Linear stability analysis
@@
Since in the MF approximation the local order is slaved by the global order, it suffices to consider the linear stability analysis for the global order. Let us consider a small perturbation of the global order around the steady-state value $\textbf{m}_{\rm s}$, i.e., $\mathbf{m}(t)=\mathbf{m}_{\rm s}+\mathbf{\delta m}(t)$. The most interesting steady-state value to consider is given by the disordered state $\mathbf{m}_{\rm s}=0$, in which case the probability (S27)-(S28) can be expanded as
@@
Note that $\hat{M}_{\mu\mu}$ can be written in terms of the steady-state solution for the local order given by Eq. (S20), i.e., $\hat{M}_{\mu\mu}=zq^{\mu\mu}_{s}-1$, which we will make use of in the remaining calculation. Furthermore, the off-diagonals are related to the steady-state of the local order between the lattices, i.e., $\hat{M}_{ab}+\hat{M}_{ba}=2q^{ab}_{s}$. Hence, a perturbation of the global order couples to the steady-state value of the local order. The solution of Eq. (S47) can be written in terms of an eigenmode expansion
@@
For fixed $z$, both equations can be solved explicitly to obtain expressions for the critical values of the parameters $J_{\mu}$ and $K_{\mu}$ on the line of Hopf bifurcations. Note that Eq. (S56) sets a direct constraint on the steady-state local order values.
@@
where the latter equation is directly satisfied for $K\neq 0$ (a proof is given in S4.2.2.). Hence, also in the “monomer” expression we find a critical value for the local order, which is given by $q_{\rm crit}\equiv 1/z$. Upon inserting Eq. (S48) into Eq. (S58), one can obtain an explicit result for the Hopf line in the $(J,K)-$plane.
@@
In the BG approximation, we can no longer neglect the local order for the linear stability analysis. Hence, we consider a small perturbation of the global and local order around their respective steady-state values
@@
S4.2.1 Steady-state local order
The trivial steady state is given by the disordered state with $m^{\mu}_{s}=0$ and $q^{ab}_{s}=0$. To solve for the steady state of the local order, denoted as $q^{aa}_{s}=q^{bb}_{s}\equiv q_{s}(J,K)$, we need to solve
@@
S4.2.2 Linear stability of global order
@@
S4.2.3 Linear stability of local order
@@
Appendix S5 Divergence of the residence time upon approaching the SNIPER bifurcation
@@
Appendix S6 Proof of existence of static global order on the finite square lattice
Here, we show that for any finite system size ($N<\infty$) there is at least a regime with $0<|K|\leq\mathcal{O}(1/N)$ where the static global order in the two-dimensional square-lattice is not destroyed. Moreover, in the thermodynamic scaling limit $N\rightarrow\infty$ and $|K|N={\rm constant}>0$ we show that there exists a regime for a nonzero magnetic field $h>0$ and for any coupling strength $J\geq J_{0}$ where the static global order must be preserved.
S6.1 Recap of the global order
@@
Let us first recall the steady-state global order for $K=0$, where we have two independent Ising systems on lattice $a$ and $b$, respectively. Let $P_{0,\rm s}(\boldsymbol{\sigma})\equiv\lim\limits_{t\rightarrow\infty}P(\boldsymbol{\sigma};t)|_{K=0}$ be the steady-state probability for $K=0$ and an arbitrary magnetic field $h$ (which we need later), and
$w_{0,i}^{\mu}(\sigma^{\mu}_{i})\equiv(1/2\tau)\left[1-\sigma^{\mu}_{i}\tanh{\left(J\sum\nolimits_{\langle i|j\rangle}\sigma^{\mu}_{j}+h\right)}\right]$ (S95)
the transition rate when $K=0$ in the presence of a magnetic field $h$. The steady-state global order for $K=0$, denoted as $m^{\mu}_{0,\rm s}$, reads [5]
$m^{\mu}_{0,\rm s}\equiv\lim\limits_{h\rightarrow 0^{\pm}}N^{-1}\sum_{\{\boldsymbol{\sigma}\}}\sum_{i=1}^{N}\sigma^{\mu}_{i}P_{0,\rm s}(\boldsymbol{\sigma}),$ (S96)
which is nonzero for $J\gtrsim\ln{(1+\sqrt{2})}/2$ when $N\gg 1$ (for $N\rightarrow\infty$ the sign $\gtrsim$ changes to $>$). Note that the limit $h\rightarrow 0^{\pm}$ of the magnetic field weakly breaks the $\mathbb{Z}_{2}$-symmetry, such that the steady-state magnetization does not correspond to the unstable value with $m^{\mu}_{0,s}=0$.
@@
$\delta w_{i}^{\mu}(\sigma^{\mu}_{i})=(1/2\tau)[1-2\delta_{\mu,b}]\sigma^{a}_{i}\sigma^{b}_{i}\operatorname{sech}^{2}{\left(J\sum\nolimits_{\langle i|j\rangle}\sigma^{\mu}_{j}+h\right)},$ (S98)
@@
$\delta m^{\mu}_{\rm s}\equiv\lim\limits_{h\rightarrow 0^{\pm}}N^{-1}\sum_{\{\boldsymbol{\sigma}\}}\sum_{i=1}^{N}\sigma^{\mu}_{i}\delta P_{\rm s}(\boldsymbol{\sigma}).$ (S100)
@@
S6.4 Bound on perturbations of global order
@@
$\delta m^{\mu}_{\rm s}$ $\leq\lim\limits_{h\rightarrow 0^{\pm}}|N^{-1}\sum_{\{\boldsymbol{\sigma}\}}\sum_{i=1}^{N}\sigma^{\mu}_{i}\delta P_{\rm s}(\boldsymbol{\sigma})|$
$\leq\lim\limits_{h\rightarrow 0^{\pm}}N^{-1}\sum_{\{\boldsymbol{\sigma}\}}\sum_{i=1}^{N}|\sigma^{\mu}_{i}\delta P_{\rm s}(\boldsymbol{\sigma})|$
$=\lim\limits_{h\rightarrow 0^{\pm}}N^{-1}\sum_{\{\boldsymbol{\sigma}\}}\sum_{i=1}^{N}|\sigma^{\mu}_{i}||\delta P_{\rm s}(\boldsymbol{\sigma})|$
$=\lim\limits_{h\rightarrow 0^{\pm}}\sum_{\{\boldsymbol{\sigma}\}}|\delta P_{\rm s}(\boldsymbol{\sigma})|.$ (S101)
@@
$\delta m^{\mu}_{\rm s}$ $\geq\lim\limits_{h\rightarrow 0^{\pm}}-|N^{-1}\sum_{\{\boldsymbol{\sigma}\}}\sum_{i=1}^{N}\sigma^{\mu}_{i}\delta P_{\rm s}(\boldsymbol{\sigma})|$
$\geq\lim\limits_{h\rightarrow 0^{\pm}}-N^{-1}\sum_{\{\boldsymbol{\sigma}\}}\sum_{i=1}^{N}|\sigma^{\mu}_{i}\delta P_{\rm s}(\boldsymbol{\sigma})|$
$=\lim\limits_{h\rightarrow 0^{\pm}}-N^{-1}\sum_{\{\boldsymbol{\sigma}\}}\sum_{i=1}^{N}|\sigma^{\mu}_{i}||\delta P_{\rm s}(\boldsymbol{\sigma})|$
$=\lim\limits_{h\rightarrow 0^{\pm}}-\sum_{\{\boldsymbol{\sigma}\}}|\delta P_{\rm s}(\boldsymbol{\sigma})|$ (S102)
@@
$|\delta m^{\mu}_{\rm s}|\leq\lim\limits_{h\rightarrow 0^{\pm}}\sum_{\{\boldsymbol{\sigma}\}}|\delta P_{\rm s}(\boldsymbol{\sigma})|.$ (S103)
This is our first main result. We are left with determining an upper bound for $\lim_{h\rightarrow 0^{\pm}}\sum_{\{\boldsymbol{\sigma}\}}|\delta P_{\rm s}(\boldsymbol{\sigma})|$, for which we use the following theorem shown in [6] (Theorem 2.1):
@@
Since it is known that the finite volume Glauber dynamics on $\mathbb{Z}^{d}$ is exponentially weakly ergodic (see Theorem 3.3 on page 117 in [7] together with proposition 3.9 on page 124 or simply use Eq. (3.15) in [7]), we can directly use Theorem 1. To translate the results from Theorem 1 into a bound for $|\delta m^{\mu}_{\rm s}|$, we first want to bound the subordinate norm for the perturbed infinitesimal generator. To do so, note that the perturbation to the transition rates obeys the following bound,
$|K\delta w_{i}^{\mu}(\sigma^{\mu}_{i})|=|(K/2\tau)[1-2\delta_{\mu,b}]\sigma^{a}_{i}\sigma^{b}_{i}\operatorname{sech}^{2}{(J\sum\nolimits_{\langle i|j\rangle}\sigma^{\mu}_{j}+h)}|\leq|K|/2\tau,$ (S106)
@@
where we used that $|\operatorname{sech}^{2}(x)|\leq 1$ for $x\in\mathbb{R}$. These perturbations enter the off-diagonal terms of $\mathbf{\hat{A}}$, so we have established that $\hat{A}_{ij}\leq|K|/2\tau$ for $i\neq j$. Since we consider single spin-flip dynamics, each row/column in $\mathbf{\hat{A}}$ has $2N$ nonzero entries excluding the diagonal entry, which is equal to $\hat{A}_{ii}=-\sum_{j}\hat{A}_{ij}$. Therefore, the diagonal term can also be bounded by $|\hat{A}_{ii}|\leq 2N\times|K|/2\tau$. Combining these results, we obtain the following bound for the subordinate norm of the perturbed infinitesimal generator,
$\|\mathbf{\hat{A}}\|\leq 2\times(2N)\times|K|/2\tau=2N|K|/\tau.$ (S107)
@@
Finally, we note that $\|\mathbf{p}_{\rm s}-\mathbf{\hat{p}}_{\rm s}\|$ in Theorem 1 translates into $\lim_{h\rightarrow 0^{\pm}}|K|\sum_{\{\boldsymbol{\sigma}\}}|\delta P_{\rm s}(\boldsymbol{\sigma})|$ in our work. Combining all together, we obtain that there exists $b>0$ and $c>2$ such that for $K\neq 0$
$\lim\limits_{h\rightarrow 0^{\pm}}\sum_{\{\boldsymbol{\sigma}\}}|\delta P_{\rm s}(\boldsymbol{\sigma})|\leq\frac{2N(1+\ln{(c/2)})}{b\tau},$ (S108)
@@
and therefore the perturbation is bounded by a finite number (when $N<\infty$ and $\tau\neq 0$) independent of $K$ up to first order
$|\delta m^{\mu}_{\rm s}|\leq\frac{2N(1+\ln{(c/2))}}{b\tau}<\infty.$ (S109)
S6.5 Connecting the dots